# Predicting the Distribution of Treatment Effects: A Covariate-Adjustment Approach

### v1

@@ 2.1 Conditions for Point Identification
Blue lines represent the distribution of $Y(1)$ conditional on a given value of $X=x$, and red lines the distribution of $Y(0)$. Panels in the first row show the cumulative distributions and the second row the probability density. Left column shows a scenario where there is partial identification, and right column a scenario with point identification. Green dashed line represents a point $t$ where $F_{1}(t|x)-F_{0}(t|x)$ is maximized. In both scenarios $\min_{t}F_{1}(t|x)-F_{0}(t|x)=0$, for example for $t=0$.
@@
From equation (2.5), a sufficient condition for point identification of $\theta$ is that $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ for all $x$. Figure 1 gives intuition for when this conditions holds, focusing on a single value of $X=x$ for ease of exposition. All panels show the conditional distributions of $Y(1)$ (in blue) and $Y(0)$ (in red), which represent the prediction of the potential outcomes given $X=x$. The upper panels show the cumulative distribution, and the lower panels the probability density. Note that predictions are not a single point since potential outcomes are random variables even conditional on $X=x$. In both scenarios $\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=0$, for example for $t=0$, and $F_{1}(t|x)-F_{0}(t|x)$ is maximized at $t=0.75$ (green dashed line).
In the lower left panel, the distributions of $Y(1)$ and $Y(0)$ are partially overlapping, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}<1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper left panel). Since there is overlap, one cannot accurately predict if $Y(1)-Y(0)<0$ or not. For example, if an individual is at the right tail of $Y(1)$ (blue) but at the left tail of $Y(0)$ (red), then $Y(1)-Y(0)>0$. If the opposite happens, then $Y(1)-Y(0)<0$. However, since the overlap is small, the probability that $Y(1)-Y(0)<0$ is large, and the fraction harmed (for this value of $X=x$) is close to point identified. In the lower right panel the distributions are disjoint, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper right panel). In this case, even though the covariates are not informative enough to get a point-prediction of $Y(1)$ and $Y(0)$, it is informative enough to learn that $Y(1)<Y(0)$ for this value of $X=x$, and thus all individuals with $X=x$ are harmed.
Note that the interval in (2.5) is an average of $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ and $1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ over all values of $X$. Hence, the figure illustrates that the informativeness of covariates in predicting the outcomes is crucial for having an narrow interval for $\theta$. In fact, if the predictions (conditional distributions) of $Y(1)$ and $Y(0)$ are disjoint for all values of $X$, then $\theta$ is point-identified. Note that in Figure 1 the green line represents the value of $s^{*}_{L}(x)$, which is the point $t$ that maximizes the distance $F_{1}(t|x)-F_{0}(t|x)$. Hence, $s^{*}_{L}(x)$ is the point that makes the two distributions as separated as possible. This fact is what motivates the notation $s$, for separator function.
@@ 3 Finite Sample Inference with Sample-Splitting
In this section, I propose sample-splitting estimators for $(\theta_{L}^{*},\theta_{U}^{*})$ and a confidence interval (CI) for $\theta$ that provides finite-sample guarantee of coverage. Given an iid sample $(Y_{i},D_{i},X_{i})_{i=1}^{n}\sim P$ that satisfies A1, the estimation strategy consists of randomly splitting the sample into two, using the first set to estimate $(s^{*}_{L},s^{*}_{U})$ with $(\hat{s}_{L},\hat{s}_{U})$, and the second set to estimate the bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ (as in 2.3) with a sample analogue. Conditional on the first sample, these bounds are valid due to (2.1), and finite sample inference is achieved by applying the DKW inequality (Dvoretzky et al., 1956; Massart, 1990).
@@
$F_{1}(t|X)$ and $F_{0}(t|X)$ can be estimated with a variety of methods, including machine learning algorithms or other nonparametric estimators (see, e.g., Kneib et al., 2023 for a review). Guidelines for estimation of $\hat{s}$ are discussed in Appendix C. Although any method can be used to estimate $F_{1}(t|X)$ and $F_{0}(t|X)$, estimates $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$ yield bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ that are closer to the sharp $[\max_{t}\Delta(t,s^{*}_{L}),1+\min_{t}\Delta(t,s^{*}_{U})]$, and poor estimators can lead to wider intervals.
A default choice is to make the main and auxiliary samples roughly the same size, but that is not necessary. In general, a larger auxiliary sample helps estimate $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$, which improves the length of the target bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$. On the other hand, a larger auxiliary sample implies a smaller main sample, which leads to larger variance of $(\widetilde{\theta}_{L},\widetilde{\theta}_{U})$ around the target bounds.
@@
The finite-sample guarantee comes from the DKW inequality (Dvoretzky et al., 1956; Massart, 1990), a concentration inequality for empirical cdfs. A similar approach was implemented in Ruiz and Padilla (2022) in the context of finite-sample inference on $\theta$ in the absence of covariates. My proof differs in that I consider the splitting of the sample, and I use the one-sided version of the DKW inequality in Massart (1990) to get a smaller critical value for the one-sided CIs.
Note that a large main sample leads to $c_{\alpha}$ being close to zero, and a large auxiliary sample makes a suitable choice of $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$. Therefore, under regularity conditions, the CIs in Theorem 2 collapse into the sharp bounds in (2.5) as both sample sizes grow to infinity. The next section investigates the large sample behavior of a similar estimator, one that allows for cross-fitting instead of sample-splitting.
@@ 4 Large Sample Inference with Cross-Fitting
    Randomly split the treatment sample into $K$ sets of same size, that is, a partition of $\{1\leq i\leq n:D_{i}=1\}$ into equal-sized folds $\{\mathcal{I}^{1}_{1},\dots,\mathcal{I}^{1}_{K}\}$. Similarly, split the control sample into $\{\mathcal{I}^{0}_{1},\dots,\mathcal{I}^{0}_{K}\}$, and denote $\mathcal{I}_{k}=\mathcal{I}^{1}_{k}\cup\mathcal{I}^{0}_{k}$;
@@
In order to derive the asymptotic properties of the cross-fitting estimators, I define new regularity conditions. The results are uniform over a set of probability functions $\mathcal{P}$. This is because the literature on inference in partially identified models argues that asymptotic results have to be uniform in order to obtain approximations that accurately represent the finite-sample behavior (e.g., Imbens and Manski, 2004; Andrews and Soares, 2010). To this end, a slight revision in notation is introduced, incorporating an additional subscript $P$ for $P\in\mathcal{P}$. Consequently, $\Delta$ in (2.4) is denoted as $\Delta_{P}$, while the sharp transformations in (2.6) and (2.7) are represented by $s^{*}_{L,P}$ and $s^{*}_{U,P}$, and sharp bounds in (2.5) are $(\theta_{L,P}^{*},\theta_{U,P}^{*})$.
@@
Assumption 2(i) restricts the possibility of point masses in the distribution of the outcome. A2(ii) simply requires the estimators $(\hat{s}_{L},\hat{s}_{U})$ to have any limit in probability, at any rate of convergence, and allows them to be inconsistent to the sharp $(s^{*}_{L,P},s^{*}_{U,P})$. A2(iii) ensures the asymptotic distribution is first-order insensitive to the choice of the optimizers in Definition 2(iii), and it is a generalization for $P\in\mathcal{P}$ of the assumption used in Fan and Park (2010) in their approach to inference on Makarov bounds without covariates. Relaxing A2(iii) is possible and is discussed in Online Appendix III. A2(iv) is used to ensure that the problem is not trivial in the sense that $Var[\sqrt{n}\widehat{\theta}_{L}]$ and $Var[\sqrt{n}\widehat{\theta}_{U}]$ do not converge to zero. This assumption is required in order to use the result from Stoye (2009) to get asymptotic validity of the two-sided CI in Theorem 4.
@@
Definitions of $\sigma^{2}_{L,P}$, $\sigma^{2}_{U,P}$ and $\sigma_{L,U,P}$ are provided in Appendix D. The cross-fitting estimators are asymptotically normal at the usual $\sqrt{n}$ rate, centered at bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$ that may be wider than the sharp $[\theta_{L}^{*},\theta_{U}^{*}]$, but which always contain $\theta$. The nonparametric rates of convergence of $\hat{s}_{L},\hat{s}_{U}$ only affect the target bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$, which always contain $\theta$. The resulting $\sqrt{n}$ rate is made possible through the splitting of the sample. Once conditioning on the estimated $(\hat{s}_{L,k},\hat{s}_{U,k})$ and under A2(iii), $(\widehat{\theta}_{L},\widehat{\theta}_{U})$ behave as sample averages. Conditions under which $(\bar{\theta}_{L,P},\bar{\theta}_{U,P})=(\theta_{L,P}^{*},\theta_{U,P}^{*})$ are given in Theorem 5.
Confidence intervals for $\theta$ follow from consistent sample-analogue estimators for $\sigma_{L}^{2}$, $\sigma_{U}^{2}$ and $\sigma_{L,U}$. Expressions are provided in Appendix D. One-sided confidence intervals follow directly from the normal approximation of Theorem 3. Since $\theta\in[0,1]$, the one-sided CIs are particularly useful for testing $\theta=0$ or $\theta=1$. For two-sided inference on $\theta$, I suggest using the confidence interval $\text{CI}^{3}_{\alpha}$ proposed in Stoye (2009), here denoted by $\widehat{CI}_{\alpha}$. A precise definition is reproduced in Online Appendix II.
@@ 5 Monte Carlo Simulations
I illustrate the results of Theorems 2 and 3 with a simulation study. I also compare the performance of the sample-splitting and cross-fitting estimators, as well as the estimator proposed by Semenova (2023) and Ji et al. (2023) (denoted SJLS for the first letter of the authors’ names). Note both estimators are equivalent in the case of the DTE, and that CAIDE by construction estimates bounds that are always (weakly) narrower than SJLS (Appendix A).
For $n\in\{500,2000\}$, I simulate $10,000$ times from the following data generating process (DGP):
@@
For this DGP, the value of $\theta$ is $\theta_{0}\approx 0.43$. I fit with $K=5$ and $\delta=0$ the models Quantile Neural Networks, Support Vector Machine (Regression) and No Covariates ($s(x)=0$). I also calculate an Oracle model that uses the sharp transformations defined in Theorem 1, which are in general unkown if the DGP is unkown. Technical details are delayed to Appendix C. I consider two scenarios: (i) $p=10$, where only covariates $X_{1}$ through $X_{10}$ are observed; (ii) $p=20$, where all $20$ covariates are observed. Note that when $p=20$, $\theta$ is point identified since potential outcomes are uniquely determined by all $20$ covariates.
Note: Green dashed lines represent the sharp identified set of $\theta$ when $p=10$, and orange dashed lines represent the true value of $\theta$. Boxplots are composed by the median, two hinges at the first and third quartiles, and whiskers extending to the most extreme data points no further than 1.5 times the interquartile range (third minus first quartile) from the hinge.
@@
Figure 2 shows the distribution of the estimators CAIDE-CF (Cross Fitting) and SJLS for the lower and upper bounds on $\theta$, for $n=500,2000$ and scenarios $p=10,20$. It illustrates the fact that the estimators are centered around bounds determined by the different models for estimating $\hat{s}$. Some models perform better than others in yielding narrower bounds. In most cases, for example, the Regr. SVM model led to the narrowest bounds, close to the sharp bounds in the $p=10$ case. A larger sample size leads to smaller variation in the estimators. Observing all $p=20$ covariates, instead of $p=10$, can lead to narrower bounds (e.g., Regr. SVM and $n=2000$) but also to wider bounds if the sample is smaller (Quant. NNet and $n=500$). Throughout, the estimates of CAIDE are narrower than SJLS.
In Table 1, I show the probability of rejecting the hypotheses $H_{0}:\theta=0$ and $H_{0}:\theta\geq\theta_{0}$, using the one-sided confidence interval based on the lower bound in Theorem 4. Here, $\theta_{0}\approx 0.43$ is the true value of $\theta$ for this DGP. A high rejection rate for the first hypothesis indicates power, since 00 is not in the sharp identified set of $\theta$. The rejection rate of the second hypothesis indicates size and ideally should be smaller than or equal to the nominal $0.05$. I also measure the average length of the confidence intervals, defined as the distance between the lower and upper bounds calculated from the one-sided CIs in Theorem 4. I compare $n=500,2000$ and $p=10,20$, for different models and estimators. CAIDE-SS is the sample-splitting estimator of Section 3, and CAIDE-CF is the cross-fitting estimator of Section 4. The table shows that including covariates, having a larger sample size, or observing all $20$ covariates all increase power while preserving size. The sample-splitting estimators are conservative, and exhibit the largest average length for Regr. SVM. Still, it often demonstrates some power, for example it excludes zero in $99.4\%$ of the cases when $p=20$, $n=2000$ and the model is Regr. SVM. The cross-fitting estimator performs better than SJLS in terms of smaller confidence intervals while preserving coverage, and it achieves the nominal coverage rate for the oracle model when $p=20$.

### v2

@@ 2.1 Conditions for Point Identification
Blue lines represent the distribution of $Y(1)$ conditional on a given value of $X=x$, and red lines the distribution of $Y(0)$. Panels in the first row show the cumulative distributions, and in the second row, the probability density functions. The left column shows a scenario with partial identification, and the right column shows a scenario with point identification. The green dashed line represents the point $t$ that maximizes $F_{1}(t|x)-F_{0}(t|x)$. In both scenarios $\min_{t}F_{1}(t|x)-F_{0}(t|x)=0$, for example for $t=0$.
@@
From equation (2.5), a sufficient condition for point identification of $\theta$ is that $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ for all $x$. Figure 1 gives intuition for when this condition holds, focusing on a single value of $X=x$ for ease of exposition. All panels show the conditional distributions of $Y(1)$ (in blue) and $Y(0)$ (in red), which represent the prediction of the potential outcomes given $X=x$. The upper panels show the cumulative distribution and the lower panels show the probability density. Note that predictions are not a single point since potential outcomes are random variables even conditional on $X=x$. In both scenarios $\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=0$, for example for $t=0$, and $F_{1}(t|x)-F_{0}(t|x)$ is maximized at $t=0.75$ (green dashed line).
In the lower left panel, the distributions of $Y(1)$ and $Y(0)$ are partially overlapping, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}<1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper left panel). Since there is overlap, one cannot accurately predict if $Y(1)-Y(0)<0$ or not. For example, if an individual is at the right tail of $Y(1)$ (blue) but at the left tail of $Y(0)$ (red), then $Y(1)-Y(0)>0$. If the opposite happens, then $Y(1)-Y(0)<0$. However, since the overlap is small, the probability that $Y(1)-Y(0)<0$ is large, and the fraction harmed (for this value of $X=x$) is close to point identified. In the lower right panel, the distributions are disjoint, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper right panel). In this case, even though the covariates are not informative enough to get a point prediction of $Y(1)$ and $Y(0)$, it is informative enough to learn that $Y(1)<Y(0)$ for this value of $X=x$, and thus all individuals with $X=x$ are harmed.
Note that the interval in (2.5) is an average of $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ and $1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ over all values of $X$. Hence, the figure illustrates that the informativeness of covariates in predicting the outcomes is crucial for having a narrow interval for $\theta$. In fact, if the predictions (conditional distributions) of $Y(1)$ and $Y(0)$ are disjoint for all values of $X$, then $\theta$ is point-identified. Note that in Figure 1, the green line represents the value of $s^{*}_{L}(x)$, which is the point $t$ that maximizes the distance $F_{1}(t|x)-F_{0}(t|x)$. Hence, $s^{*}_{L}(x)$ is the point that makes the two distributions as separated as possible. This fact is what motivates the notation $s$, for separator function.
@@ 3 Finite Sample Inference with Sample-Splitting
In this section, I propose sample-splitting estimators for $(\theta_{L}^{*},\theta_{U}^{*})$ and a confidence interval (CI) for $\theta$ that provides a finite sample guarantee of coverage. Given an iid sample $(Y_{i},D_{i},X_{i})_{i=1}^{n}\sim P$ that satisfies A1, the estimation strategy consists of randomly splitting the sample into two, using the first set to estimate $(s^{*}_{L},s^{*}_{U})$ with $(\hat{s}_{L},\hat{s}_{U})$, and the second set to estimate the bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ (as in 2.3) with a sample analogue. Conditional on the first sample, these bounds are valid due to (2.1), and finite sample inference is achieved by applying the DKW inequality (Dvoretzky et al., 1956; Massart, 1990).
@@
$F_{1}(t|X)$ and $F_{0}(t|X)$ can be estimated with a variety of methods, including machine learning algorithms or other nonparametric estimators (see, e.g., Kneib et al., 2023 for a review). I discuss guidelines for estimating $\hat{s}$ in Appendix A.3. Although any method can be used to estimate $F_{1}(t|X)$ and $F_{0}(t|X)$, estimates $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$ yield bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ that are closer to the sharp $[\max_{t}\Delta(t,s^{*}_{L}),1+\min_{t}\Delta(t,s^{*}_{U})]$, and poor estimators can lead to wider intervals.
A default choice is to make the main and auxiliary samples roughly the same size, but that is unnecessary. In general, a larger auxiliary sample helps estimate $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$, which improves the length of the target bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$. On the other hand, a larger auxiliary sample implies a smaller main sample, which leads to larger variance of $(\widetilde{\theta}_{L},\widetilde{\theta}_{U})$ around the target bounds.
@@
The finite-sample guarantee comes from the DKW inequality (Dvoretzky et al., 1956; Massart, 1990), a concentration inequality for empirical cdfs. A similar approach was implemented in Ruiz and Padilla (2022) in the context of finite-sample inference on $\theta$ without covariates. My proof differs in that I consider the splitting of the sample, and I use the one-sided version of the DKW inequality in Massart (1990) to get a smaller critical value for the one-sided CIs.
Note that a large main sample leads to $c_{\alpha}$ being close to zero, and a large auxiliary sample makes a suitable choice of $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$. Therefore, under regularity conditions, the CIs in Theorem 2 collapse into the sharp bounds in (2.5) as both sample sizes grow to infinity. The next section investigates the large sample behavior of a similar estimator, which allows for cross-fitting instead of sample-splitting.
@@ 4 Large Sample Inference with Cross-Fitting
    Randomly split the treatment sample into $K$ sets of the same size, that is, a partition of $\{1\leq i\leq n:D_{i}=1\}$ into equal-sized folds $\{\mathcal{I}^{1}_{1},\dots,\mathcal{I}^{1}_{K}\}$. Similarly, split the control sample into $\{\mathcal{I}^{0}_{1},\dots,\mathcal{I}^{0}_{K}\}$, and denote $\mathcal{I}_{k}=\mathcal{I}^{1}_{k}\cup\mathcal{I}^{0}_{k}$;
@@
In order to derive the asymptotic properties of the cross-fitting estimators, I define new regularity conditions. The results are uniform over a set of probability functions $\mathcal{P}$. This is because the literature on inference in partially identified models argues that asymptotic results have to be uniform in order to obtain approximations that accurately represent the finite-sample behavior (e.g., Imbens and Manski, 2004; Andrews and Soares, 2010). To this end, I introduce a slight revision in notation, incorporating an additional subscript $P$ for $P\in\mathcal{P}$. Consequently, $\Delta$ in (2.4) is denoted as $\Delta_{P}$, while the sharp transformations in (2.6) and (2.7) are represented by $s^{*}_{L,P}$ and $s^{*}_{U,P}$, and sharp bounds in (2.5) are $(\theta_{L,P}^{*},\theta_{U,P}^{*})$.
@@
A2(i) restricts the possibility of point masses in the distribution of the outcome. A2(ii) simply requires the estimators $(\hat{s}_{L},\hat{s}_{U})$ to have any limit in probability, at any rate of convergence, and allows them to be inconsistent for the sharp $(s^{*}_{L,P},s^{*}_{U,P})$. A2(iii) is assumed to simplify exposition. It ensures the asymptotic distribution is first-order insensitive to the choice of the optimizers in Definition 2(iii), and it is a generalization for $P\in\mathcal{P}$ of the assumption used in Fan and Park (2010) in their approach to inference on Makarov bounds without covariates. I discuss two approaches for relaxing A2(iii) in the presence of multiple optimizers in the Online Appendix (OA.3). One of them consists in using the bootstrap for inference instead of relying on normal approximations to $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$. Example 1 illustrates a data generating process where A2(iii) holds. A2(iv) is used to ensure that the problem is not trivial in the sense that $Var[\sqrt{n}\widehat{\theta}_{L}]$ and $Var[\sqrt{n}\widehat{\theta}_{U}]$ do not converge to zero. It is only used to apply the result from Stoye (2009) to get asymptotic validity of the two-sided CI in Theorem 4.
Example 1 (Uniqueness of optimizers).
Consider a data generating process where $Y(0)=f(X)+\varepsilon$, $Y(1)=Y(0)+\tau$, $\varepsilon,\tau\perp\!\!\!\perp X$ and $\mathbb{E}[Y(1)]=\mathbb{E}[Y(0)]=0$. For simplicity, I focus on a single probability function $P$, that is, $\mathcal{P}$ is a singleton. When the distributions of $Y(1)$ and $Y(0)$ are continuous and symmetric around zero, the optimal covariate-adjustment term is the same for both the lower and upper bounds and is given by $s^{*}(x)=\mathbb{E}[Y(0)|X=x]=f(x)$. The distributions of $Y(1)-s^{*}(X)$ (in blue) and $Y(0)-s^{*}(X)$ (in red) are shown in Figure 2 below. The difference $\Delta_{P}(t,s^{*})=P(Y(1)-s^{*}(X)\leq t)-P(Y(0)-s^{*}(X)\leq t)$ is uniquely maximized and minimized respectively at the green and yellow vertical dashed lines, where the probability density functions intersect.
Visualization of Example 1. Blue lines represent the distribution of $Y(1)-s^{*}(X)$, and red lines the distribution of $Y(0)-s^{*}(X)$. The left panel shows the cumulative distribution function and the right panel shows the probability density function. The green dashed line represents the unique point $t$ where $\Delta_{P}(t,s^{*})$ is maximized, and the solid green line in the left panel represents the magnitude of $\Delta_{P}(t,s^{*})$ at this point. Similarly, the yellow dashed and solid lines show the unique point that minimizes $\Delta_{P}(t,s^{*})$ and its magnitude.
Figure 2: Example of data generating process that satisfies A2(iii).
@@
Definitions of $\sigma^{2}_{L,P}$, $\sigma^{2}_{U,P}$ and $\sigma_{L,U,P}$ are provided in Appendix A.4. The cross-fitting estimators are asymptotically normal at the usual $\sqrt{n}$ rate, centered at bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$ that may be wider than the sharp $[\theta_{L}^{*},\theta_{U}^{*}]$, but which always contain $\theta$. The nonparametric rates of convergence of $\hat{s}_{L},\hat{s}_{U}$ only affect the target bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$, which always contain $\theta$. The resulting $\sqrt{n}$ rate is made possible through the splitting of the sample. Once conditioning on the estimated $(\hat{s}_{L,k},\hat{s}_{U,k})$ and under A2(iii), $(\widehat{\theta}_{L},\widehat{\theta}_{U})$ behave as sample averages. Conditions under which $(\bar{\theta}_{L,P},\bar{\theta}_{U,P})=(\theta_{L,P}^{*},\theta_{U,P}^{*})$ are given in Theorem 5.
Confidence intervals for $\theta$ follow from consistent sample-analogue estimators for $\sigma_{L}^{2}$, $\sigma_{U}^{2}$ and $\sigma_{L,U}$. I provide expressions in Appendix A.4. One-sided confidence intervals follow directly from the normal approximation of Theorem 3. Since $\theta\in[0,1]$, the one-sided CIs are particularly useful for testing $\theta=0$ or $\theta=1$. For two-sided inference on $\theta$, I suggest using the confidence interval $\text{CI}^{3}_{\alpha}$ proposed in Stoye (2009), here denoted by $\widehat{CI}_{\alpha}$. A precise definition is reproduced in Online Appendix OA.2.
@@ 5 Monte Carlo Simulations
I illustrate the results of Theorems 2 and 3 with a simulation study. I also compare the performance of the sample-splitting and cross-fitting estimators, as well as the estimator proposed by Semenova (2023) and Ji et al. (2023) (denoted SJLS for the first letter of the authors’ names). Note that both estimators are equivalent in the case of the DTE and that CAIDE by construction estimates bounds that are always (weakly) narrower than SJLS (Appendix A.1).
For $n\in\{500,2000\}$, I simulate $10,000$ times from the following data-generating process (DGP):
@@
For this DGP, $\theta=\theta_{0}\approx 0.43$. I fit with $K=5$ and $\delta=0$ the models Quantile Neural Networks, Support Vector Machine (Regression), and No Covariates ($s(x)=0$). I also calculate an Oracle model that uses the sharp transformations defined in Theorem 1, which are generally unknown if the DGP is unknown. Technical details are delayed to Appendix A.3. I consider two scenarios: (i) $p=10$, where only covariates $X_{1}$ through $X_{10}$ are observed; (ii) $p=20$, where all $20$ covariates are observed. Note that when $p=20$, $\theta$ is point identified since potential outcomes are uniquely determined by all $20$ covariates.
Note: Green dashed lines represent the sharp identified set of $\theta$ when $p=10$, and orange dashed lines represent the true value of $\theta$. Boxplots comprise the median, two hinges at the first and third quartiles, and whiskers extending to the most extreme data points no further than 1.5 times the interquartile range (third minus first quartile) from the hinge.
@@
Figure 3 shows the distribution of the estimators CAIDE-CF (Cross Fitting) and SJLS for the lower and upper bounds on $\theta$, for $n=500,2000$ and scenarios $p=10,20$. It illustrates that the estimators are centered around bounds determined by the different models for estimating $\hat{s}$. Some models perform better than others in yielding narrower bounds. For example, in most cases, the Regr. SVM model led to the narrowest bounds, close to the sharp bounds in the $p=10$ case. A larger sample size leads to smaller variation in the estimators. Observing all $p=20$ covariates, instead of $p=10$, can lead to narrower bounds (e.g., Regr. SVM and $n=2000$) but also to wider bounds if the sample is smaller (Quant. NNet and $n=500$). Throughout, the estimates of CAIDE are narrower than those of SJLS.
In Table 1, I show the probability of rejecting the hypotheses $H_{0}:\theta=0$ and $H_{0}:\theta\geq\theta_{0}$, using the one-sided confidence interval based on the lower bound in Theorem 4. Here, $\theta_{0}\approx 0.43$ is the true value of $\theta$ for this DGP. A high rejection rate for the first hypothesis indicates power since 00 is not in the sharp identified set of $\theta$. The second hypothesis’s rejection rate indicates size, and it should ideally be less than or equal to the nominal $0.05$. I also measure the average length of the confidence intervals, defined as the distance between the lower and upper bounds calculated from the one-sided CIs in Theorem 4. I compare $n=500,2000$ and $p=10,20$ for different models and estimators. CAIDE-SS is the sample-splitting estimator of Section 3, and CAIDE-CF is the cross-fitting estimator of Section 4. The table shows that including covariates, having a larger sample size, or observing all $20$ covariates increases power while preserving size. The sample-splitting estimators are conservative and exhibit the largest average length for Regr. SVM. Still, it often demonstrates some power. For example, it excludes zero in $99.4\%$ of the cases when $p=20$, $n=2000$, and the model is Regr. SVM. The cross-fitting estimator performs better than SJLS in terms of smaller confidence intervals while preserving coverage. It achieves the nominal coverage rate for the oracle model when $p=20$.